OpenStudy (anonymous):

In triangle ABC, a = 3, b = 5, and c = 7. Find the approximate value of angle A. guys i need help with this to pass because if i dont, i dont graduate. please help me

OpenStudy (trojanpoem):

 From the cosine rule: a^2 = c^2 + b^2 - 2cbcos(A) cosA = c^2 + b^2 - a^2/ 2cb Can you continue on your own ?